OpenStudy (anonymous):

What number must you add to complete the square? x2 + 8x = 9

OpenStudy (anonymous):

Answer choices: A. 4 B. 8 C. 32 D. 16

OpenStudy (anonymous):

@chycora @hugsnotughs

OpenStudy (er.mohd.amir):

x2+8x=9 x2+2*x*4+16=9+16 (x+4)^2=25 so add 16. option D.